OpenStudy (anonymous):

2*(3x+1)-3*(2x+2)=-4

OpenStudy (amtran_bus):

A quick way to check your work would be to plug each "part" or "piece" into a graphing calculator and find the zeros/interceptions.

OpenStudy (anonymous):

2(1 + 3x) + -3(2x + 2) = -4 (1 * 2 + 3x * 2) + -3(2x + 2) = -4 (2 + 6x) + -3(2x + 2) = -4 2 + 6x + -3(2 + 2x) = -4 2 + 6x + (2 * -3 + 2x * -3) = -4 2 + 6x + (-6 + -6x) = -4 2 + -6 + 6x + -6x = -4 Combine like terms 6x + -6x = 0 -4 + 0 = -4 -4 = -4 Add '4' to each side of the equation I think you can figure out the rest.